The length of a rectangle is equal to three times its width. If the perimeter is equal to 240 feet, what is the length of the recatngle? Ican only get L(3w)= 240
Answer by RAY100(1637)

Perimeter of rectangle is 2*length+2*width
but l=3w
p=2l+2w=2(3w)+2w=8w,,,but given =240
therefore w=240/8, =30
and l=3w=3*30=90
check
Per=2(30)+2(90)=60+180=240,,,,ok